OpenStudy (anonymous):

IIrma needs to rent a car for a business trip from Phoenix to San Jose (720 miles one way). She will drive round trip and will be gone a total of 8 days. The car rental company charges $ 65.00 per day and allows for 1,600 miles for Irmas's trip. There will be an additional charge of $ 0.24 per mile over the 1,600 miles allowed. How many total miles can she drive without exceeding her budget of $ 600.00?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

Well, we know that if it is $65.00 per day, and she is gone 8 days, she is spending a minimum of $520.00, leaving her $80.00 remaining for extra mileage. $80.00/$0.24=333. So she can drive a total of 1933 miles.

OpenStudy (anonymous):

Any other questions?

OpenStudy (tkhunny):

Slightly more formal and organized. Daily + First 1600 + Excess = 600 $8 * 65 + $0 * 1600 + $0.24 * E = $600 $520 + $0.24 * E = $600 $0.24 * E = $80 E = 80/0.24 = 333 and change